### v2

# Human Perception-Guided Pretraining for Increased Generalization

Abstract
Incorporating human perception into training of convolutional neural networks (CNN) has boosted generalization capabilities of such models in open-set recognition tasks. One of the active research questions is where (in the model architecture) and how to efficiently incorporate always-limited human perceptual data into training strategies of models. In this paper, we introduce MENTOR (huMan pErceptioN-guided preTraining fOr increased geneRalization), which addresses this question through two unique rounds of training the CNNs tasked with open-set anomaly detection. First, we train an autoencoder to learn human saliency maps given an input image, without class labels. The autoencoder is thus tasked with discovering domain-specific salient features which mimic human perception. Second, we remove the decoder part, add a classification layer on top of the encoder, and fine-tune this new model conventionally. We show that MENTOR’s benefits are twofold: (a) significant accuracy boost in anomaly detection tasks (in this paper demonstrated for detection of unknown iris presentation attacks, synthetically-generated faces, and anomalies in chest X-ray images), compared to models utilizing conventional transfer learning (e.g., sourcing the weights from ImageNet-pretrained models) as well as to models trained with the state-of-the-art approach incorporating human perception guidance into loss functions, and (b) an increase in the efficiency of model training, requiring fewer epochs to converge compared to state-of-the-art training methods.
1 Introduction
Figure 1: MENTOR approach. Step 1: An autoencoder-based model is first trained to recreate human saliency maps, and thus to build an understanding of human perception-sourced salient features into the encoder $\mathcal{F}_{\text{encoder}}$. Step 2: Such pre-trained encoder $\mathcal{F}_{\text{encoder}}$ is then decoupled from the autoencoder, and along with a classifier $\mathcal{F}_{\text{class}}$ are tuned for the anomaly detection task utilizing standard cross-entropy loss.
Human perceptual information is often incorporated into deep learning training strategies in order to improve generalization [3, 33], align models with human-sourced saliency [5, 10], and reduce training time by supplying additional prior knowledge into the training process [5, 43]. However, a common and valid criticism is the high cost of acquiring human annotations or eye tracking-sourced saliency data. For anomaly detection tasks, such as biometric presentation attack detection (PAD), new attack types are developed frequently. Medical diagnostics may require continual update of human perceptual understanding of anomalies present in medical samples. In all these tasks, requiring few-shot learning solutions functioning well in an open-set classification regime, collecting more training samples and/or more human saliency data is difficult or even impossible. This is why an effective use of the existing, although always-limited human saliency information originating from domain experts is crucial. An associated question is how and where to appropriately incorporate human perceptual understanding of a task into model training to deliver strong priors allowing for better generalization of such models.
This paper introduces MENTOR, a framework for human perception-guided pretraining, evaluated in the context of open-set iris PAD, synthetic face detection, and disease classification from chest X-ray scans. MENTOR relies on the intuition that (as children) we first learn the representations of the visual world without knowing the “class labels,” and we assign meaning in what we see later in our lives. Around being one year old, infants began to point, imitate, and understand objects merely by observation or visual salience, without any explicit knowledge of the object’s application or type (classification) [16]. We emulate this iterative process by first training an autoencoder to learn the associations between an input image and a human saliency map (only visual salience, with no class labels). Once this representation is learned, the decoder-side is removed, a classifier is put on top of the encoder’s latent space, and the entire model is then fine-tuned in a regular way using cross-entropy loss to solve a classification task at hand. Fig. 1 illustrates this pipeline. Our experiments show that such class label-agnostic, but human perception-aware pretraining of the model’s backbone allows for (a) better generalization to unknown samples than models fine-tuned in the same way but initialized with weights obtained in pre-training on a massive amount data such as ImageNet, (b) same as (a) when compared to models trained with a state-of-the-art human perception guidance incorporated into a loss function, and (c) faster convergence compared to models initialized with ImageNet weights. We make these observations for three different domains, in which humans can deliver meaningful saliency information (iris PAD, synthetic face detection, and chest X-ray-based diagnosis), and for three different neural network architectures (ResNet, Inception and EfficientNet). In other words, MENTOR efficiently incorporates the limited human perceptual understanding of a domain into training of convolutional neural networks (CNN).
Two core differences between MENTOR and state-of-the-art unsupervised representation learning approaches, such as DINO [6], are that (a) MENTOR leverages human perception-based guidance in the pre-training phase, and (b) MENTOR operates in a very data-limited regime, since it’s targeted to domains, in which collecting data is intricate, but humans competent in a given domain can be found. It is also noteworthy that MENTOR does not assume any particular CNN architecture, and works without any architectural changes. For clarity, this paper is organized around the following two research questions:
  • RQ1: Does MENTOR improve the performance of iris presentation attack detection, synthetic face detection, and chest X-ray-based diagnosis compared to models pre-trained with massive amount of non-human-perception-sourced data (ImageNet [13])? (answered in Sec. 5.1)
  • RQ2: Does MENTOR allow for faster model’s fine-tuning, compared to pre-training with large non-human-perception-sourced data? (answered in Sec. 5.2)
The sources codes of the proposed approach are offered with this paper [^1] Codes will be released when the peer-reviewed version is published..
2 Related Work
2.1 Human Saliency-Guided Training
Human perceptual information related to visual tasks is usually collected via image annotations [3, 5], eye tracking [4, 10], or measuring reaction times [18]. Human salience, estimated from the perceptual data, has been successfully integrated into the training process by perception-based training data augmentations [3], integrating the perceptual information into the loss function [5, 18, 36], or as a regularization approach [14]. Other approaches include using attention mechanisms, which typically require changes to the model and are architecture-specific [32]. Human-guided models have shown to improve performance [5], model interpretability [32] and explainability [9].
The most architecture-agnostic implementations of incorporating human saliency into model’s training are those using perception-specific loss function components. Boyd et al.[5] introduced the CYBORG loss function, which penalizes the model for divergence between human saliency maps and model’s Class Activation Maps (CAM) [47]. While this method aligns the activations of the model’s last convolutional layer with human saliency, it’s unclear if that mechanism actually builds a human perceptual intelligence into the model, especially in its earlier layers, or acts as a regularizer by lowering the entropy of model’s salience [36]. The MENTOR approach proposed in this paper is different from CYBORG (and other similar human perception-based training strategies) in a sense that it encourages the model to first build associations between input images and human-sourced salient features, without telling the model what task these features are useful for, to create more general interpretations of the visual world.
Sonsbeek et al.utilize a two-stage training strategy to learn global and local features [43]. Their model is first trained on a larger dataset with no human salience, then fine-tuned on a smaller dataset with human saliency maps using a knowledge distillation module, with freezing selected weights. MENTOR deploys the reverse strategy, leveraging human saliency maps first for learning class-agnostic human-sourced features, and in its second stage utilizing a regular cross-entropy-based training without a requirement to freeze any weights.
2.2 Efficient Use of Human Annotations
Crum et al. [8] applied human salience in a Teacher-Student training paradigm to make a better use of limited human salience data. In their work, human annotations were first used to train “teacher” models, which are then used to generate subsequent model salience for training “student” models. Teachers were trained using the CYBORG loss, and afterwards used to generate saliency maps on un-annotated training samples using CAM [47] and RISE [34]. This training paradigm boosted performance in synthetic face detection and iris PAD. Interestingly, the MENTOR’s byproduct is an autoencoder that generates human saliency maps for unseen data, and thus it can complement approaches such as the above teacher-student learning paradigms.
3 MENTOR Approach
3.1 Methodology
Figure 2: A byproduct of the MENTOR pre-training approach is the autoencoder $\mathcal{F}_{\text{decoder}}\big{(}\mathcal{F}_{\text{encoder}}(\cdot)\big{)}$ predicting saliency maps that resemble human salience (top row: iris presentation attack detection, middle row: synthetic face detection, and bottom row: chest X-ray-based diagnosis).
MENTOR is a novel approach of incorporating human salience into model training through a two-part training series. First, an autoencoder is trained to generate human-like saliency given an input image, but without any class labels (see “Step 1” in Fig. 1). In that way, the model creates associations between input samples and human-salient regions. Once these associations are made, we put a single, fully-connected layer (initialized with random weights) on top of the encoder’s embeddings and fine-tune the entire architecture to solve the classification task (see “Step 2” in Fig. 1).
It’s analogous to theories involving how humans gain perceptual understanding of the visual world, by first being exposed to visual stimuli without being given explanations (or “class labels”) of what they see. As shown later in Sec. 5), such human perception-based pre-training results in a better performance in three anomaly detection tasks (from three different domains) compared to those observed for models initialized with weights obtained after training with a large visual dataset (ImageNet), and than for models trained with state-of-the-art loss function-based human perception guidance.
More formally, let’s consider a state-of-the-art family of approaches, such as [5, 20, 35], incorporating human salience into loss functions used to train a model $\mathcal{F}$ in a supervised manner in one phase. Re-phrasing the CYBORG loss [5] (helpful in stressing the differences between the above approaches and MENTOR), we can summarize human-guided loss functions as:
$\mathcal{L}=\alpha\mathcal{L}_{\text{classification}}+(1-\alpha)\mathcal{L}_{\text{human-salience}}$ (1)
where $\alpha$ is a trade-off parameter weighting human- and model-based saliencies (not needed in MENTOR, as seen later), while
$\mathcal{L}_{\text{classification}}=\frac{1}{K}\sum_{k=1}^{K}\sum_{c=1}^{C}\bm{1}_{y_{k}\in\mathcal{C}_{c}}\Big{(}\log p_{\mathcal{F}}(y_{k}\in\mathcal{C}_{c})\Big{)}$ (2)
is a cross-entropy-based classification loss function, and
$\mathcal{L}_{\text{human-salience}}=\frac{1}{K}\sum_{k=1}^{K}\sum_{c=1}^{C}\mathcal{D}\Big{(}\textbf{s}_{k,\mathcal{C}_{c}}^{\text{human}}-\textbf{s}_{k,\mathcal{C}_{c}}^{\mathcal{F}}\Big{)}$ (3)
where $y_{k}$ is a class label for the $k$-th sample, $\bm{1}$ is an indicator function equal to $1$ when $y_{k}\in\mathcal{C}_{c}$, and 0 otherwise, $C$ is the total number of classes, $K$ is the number of samples in a batch, $\textbf{s}_{k,\mathcal{C}_{c}}^{\text{human}}$ and $\textbf{s}_{k,\mathcal{C}_{c}}^{\mathcal{F}}$ are the human and the model $\mathcal{F}$’s saliency for the $k$-th sample representing class $\mathcal{C}_{c}$, respectively, and finally $\mathcal{D}$ is the measure of dissimilarity between saliency maps (e.g.,the Mean Square Error distance).
MENTOR, in contrast to the above formulation, disentangles the human-perception-guided pre-training and classification fine-tuning of the model $\mathcal{F}$ consisting of two components: the encoder part $\mathcal{F}_{\text{encoder}}$, learning human salience, and classification part $\mathcal{F}_{\text{class}}$. Namely, in the first step (cf. Fig. 1) we use the human saliency to train $\mathcal{F}_{\text{encoder}}$ using the following $\mathcal{L}_{\text{human-salience}}^{\text{MENTOR}}$ loss:
$\mathcal{L}_{\text{human-salience}}^{\text{MENTOR}}=\frac{1}{K}\sum_{k=1}^{K}\big{\|}\textbf{s}_{k}^{\text{human}}-\textbf{s}_{k}^{\mathcal{F}_{\text{decoder}}(\mathcal{F}_{\text{encoder}})}\big{\|}^{2}$ (4)
[…]
Note that the decoder part $\mathcal{F}_{\text{decoder}}$ is not utilized in Step 2, however the entire autoencoder trained in Step 1 can also be used to generate human-like salience maps and replace real salience maps in previosly-proposed human saliency-based training paradigms, such as [5, 8, 20] or [35].
3.2 Neural Network Architectural Choices
The MENTOR approach does not specify a type of autoencoder. We experimented with UNET [38] and UNET++ [48] with three different backbones: ResNet152 [17], Inception-V4 [39] and EfficientNet-b7 [40]) using the same training configurations [19] to observe stability of the proposed approach across visual domains and architectures. Specific domains and backbones for the autoencoder reported in Sec. 5 are as follows: UNET (all backbones for iris PAD domain) and UNET++ (all backbones for synthetic face detection and chest X-ray-based diagnosis).
Also, MENTOR does not specify the classifier put in Step 2 on top of the encoder embeddings. We demonstrate effectiveness of this approach by adding a single-layer (linear) classifier to minimize the extra architectural components added to human saliency-guided pre-trained encoder, but there are no theoretical reasons for not exploring the generalization gain achieved for deeper structures as a future research topic.
3.3 Training and Evaluation Details
MENTOR does not need any specific optimizer or hyperparameter settings. For reference purposes, we provide settings we have used in this work. The autoencoders in Step 1 were trained with the AdamW optimizer, with a learning rate of 0.0001, and with a batch size of 8. MSE loss was used to assess the agreement bewteen human saliency and the predicted saliency maps, all scaled to a canonical $224\times 224$ pixel resolution. Training was continued until 50 epochs, but many models converged quickly (in less than 5 epochs). All full models in Step 2 were trained using cross-entropy loss using Stochastic Gradient Descent (SGD), with a learning rate of 0.005 decreased by 0.1 every 12 epochs. As in Step 1, maximum number of epochs was 50.
We evaluate the proposed MENTOR approach using Area Under the Receiver Operating Characteristic Curve (AUROC). For all experiments we perform ten independent training runs instantiated using different seeds. We use boxplots with notches representing 95% confidence intervals of the median values for visual assessment of the observed differences between methods. For formal assessment of statistical significance of these differences, we use two-sample one-tailed Kolmogorov-Smirnov test with the null hypothesis stating the equality of mean AUROCs, and the alternative hypothesis stating better AUROC achieved by MENTOR. In addition, we also report the performance of the best three performing models (out of the ten training runs). The latter helps in estimating the highest performance that can be achieved when “cherry-picking” the models is justified (e.g.,for final deployment).
4 Datasets
4.1 Training and Validation Data Subsets
Domain 1: Iris Presentation Attack Detection
For training the autoencoder (“Step 1” in Fig. 1), we use the only (known to us) dataset of bona fide and anomalous iris images accompanied by human saliency maps, collected by Boyd et al.[3]. The bona fide and abnormal iris images for this experiment were sampled from a superset of already published live iris and iris presentation attack sets [1, 31, 15, 29, 46, 42, 30, 44, 41, 45, 12]. In Boyd et al.’s experiments, participants recruited via Mechanical Turk were asked whether the iris was bona fide or abnormal, and were instructed to hand-annotate regions of the image which support their decision. Only correctly classified samples were used in post-processing. Annotations of the same image originating from multiple subjects were averaged together, resulting in 765 saliency maps (see Fig. 2 for examples). These 765 images and saliency maps were randomly divided by the authors into train (n=612) and validation (n=153) sets, and this split was used in training of all the autoencoder instances in this work.
For training the subsequent iris PAD classification models (“Step 2” in Fig. 1), the same 765 images (but now with class labels and no salience maps) were used as training images, in conjunction with additional 23,312 validation images sampled from the same superset as used in [3]. The train and validation sets are subject-disjoint.
Domain 2: Synthetic Face Detection
We trained the autoencoder using real and synthetically-generated face images and human saliency maps offered by [5]. 363 humans recruited via Amazon Mechanical Turk annotated regions of the image deemed important to their classification decision (real / fake). Only correctly classified samples (images and accompanying saliency maps) were used during training. For simplicity, 765 samples were randomly selected without replacement to match the number of samples available for Domain 1 (Iris Presentation Attack Detection).
For training the subsequent synthetic face classification models (“Step 2” in Fig. 1), all 765 images were used as training images in conjunction with additional 20,000 validation samples extracted from the same sources as in [5]. Train and validation sets were subject-disjoint.
Domain 3: Chest X-ray-based Diagnosis
The autoencoder was trained with normal and abnormal chest X-ray images [21], and corresponding eye-tracking visual salience collected from radiologists examining these X-ray scans [2]. We consider only abnormal categories represented by at least 200 training images, that is: atelectasis, cardiomegaly, edema, lung opacity, pleural effusion, pneumonia, and support devices. We removed images containing erroneous class labels and consider only images where radiologists were certain of their classification decision. The intensity of pixels in eye tracking-based salience maps is proportional to the strength of the eye fixation. If a sample included human saliency maps from multiple radiologists, we combined these maps by taking the maximum intensity for each pixel. This was done to preserve salient information provided by each expert, who may consider different features when classifying the samples, instead of considering only the most “popular” features. All images were resized to $224\times 224$. For consistency, 765 samples and accompanying saliency maps were used during the training phases to match with sample sizes used in two other domains.
For training the subsequent chest X-ray classification models (“Step 2” in Fig. 1), all 765 images were used as training images, in conjunction with additional 4,878 subject-disjoint validation samples sampled from the same dataset.
4.2 Test Data Sets
Domain 1: Iris Presentation Attack Detection
We evaluate MENTOR using the Iris Liveness Detection 2020 Competition (LivDet-2020) official test set [11]. LivDet-2020 is a competition held to benchmark iris presentation attack algorithms, and 2020 edition included the largest number of abnormal (from iris recognition point of view) classes, including: artificial eyeballs with irises printed on them, textured contact lenses, postmortem iris images, paper print outs, synthetically-generated iris samples, images of diseased eyes, and images of eyes wearing textured contact lenses printed on paper and then re-photographed. The evaluation made with the LivDet-Iris 2020 test data is thus quite rigorous as it tests the algorithms in an open-set scenario, in which attack types unseen during training are used in testing.
Domain 2: Synthetic Face Detection
We evaluate MENTOR models detecting synthetic faces using (a) synthesized images sampled from six different GAN architectures (ProGAN [22], StarGANv2 [7], StyleGAN [23], StyleGAN2 [28], StyleGAN2-ADA [25], and StyleGAN3 [26], and (b) authentic face images from FFHQ [27] and CelebA-HQ [24], comprising of 7,000 test images. These GAN models and the source of live faces were not considered in generating the train and validation sets, hence the open-set regime of this evaluation protocol is kept. Future works may include testing with additional synthetic face image benchmarks, incorporating for instance stable diffusion-based samples [37], if such formal benchmarks are available.
Domain 3: Chest X-ray-based Diagnosis
The test set was comprised of 8,265 samples extracted from [21], completely disjoint from previous training splits. Samples with at least one abnormal category (atelectasis, cardiomegaly, edema, lung opacity, pleural effusion, pneumonia, or support devices) were classified as abnormal, with the contrary being normal samples.
5 Results
To answer both research questions, we compare MENTOR-trained models with those (a) pre-trained using a classical transfer learning approach and initialized with weights resulting from training the models with ImageNet samples, and (b) trained with a state-of-the-art human perception-guided approach (CYBORG) after initializing the weights with ImageNet-trained models’ weights. The graphical summaries (in a form of boxplots, along with numerical values for mean and standard deviations) of the accuracy for each domain-architecture combination are shown in Figs 3-5. The extra top-three-models-based results are summarized in Tab. 1. Validation accuracies as a function of epoch number for all domain-architecture combinations are shown in Figs 6-8.
Table 1: Mean and standard deviations of the top three AUROC scores (calculated independently for each combination of domain-architecture).
Dataset ImageNet ImageNet MENTOR
       Backbone + Xent + CYBORG [5] (this paper)
Iris PAD
       ResNet152 0.900$\pm$0.006 0.914$\pm$0.003 0.939$\pm$0.009
       Inception-V4 0.853$\pm$0.029 0.830$\pm$0.016 0.939$\pm$0.012
       EfficientNet-7b 0.895$\pm$0.005 0.832$\pm$0.007 0.922$\pm$0.004
Synthetic Face Detection
       ResNet152 0.579$\pm$0.007 0.568$\pm$0.045 0.601$\pm$0.011
       Inception-V4 0.766$\pm$0.016 0.743$\pm$0.005 0.789$\pm$0.006
       EfficientNet-7b 0.489$\pm$0.009 0.535$\pm$0.018 0.600$\pm$0.024
Chest X-ray-based diagnosis
       ResNet152 0.852$\pm$0.001 0.864$\pm$0.002 0.864$\pm$0.002
       Inception-V4 0.848$\pm$0.002 0.846$\pm$0.003 0.856$\pm$0.001
       EfficientNet-7b 0.827$\pm$0.002 0.827$\pm$0.002 0.845$\pm$0.005
5.1 Answering RQ1 (improvement of generalization)
MENTOR pre-training compares very favorably to models initialized with ImageNet weights, as well as to models trained with a state-of-the-art human-perception training paradigm (CYBORG) using the same data and the same human saliency maps. In six out of nine domain-architecture combinations, MENTOR achieves statistically significantly better average AUROC (across 10 independent training runs) than CYBORG- and pure cross-entropy-based training (cf. Figs 3-5).
Looking only at top-three average AUROC scores (Tab. 1), MENTOR pre-training outperforms both models initialized with ImageNet-sourced weights and (a) fine-tuned on domain-specific data with cross-entropy loss (without human-saliency), and (b) fine-tuned using CYBORG loss (i.e.,with human-saliency).
Thus, answering RQ1, we conclude that MENTOR pre-training is an effective method to improve the generalization capabilities in iris PAD, synthetic face detection, and chest x-ray-based diagnosis tasks.
5.2 Answering RQ2 (fine-tuning effectiveness)
To answer RQ2, we qualitatively evaluated model’s validation accuracy during training for all combinations of the domain and network architecture. As seen in Figs. 6-8, MENTOR helps reduce training time by bringing the models to the convergence point faster. We observe higher validation accuracy and smaller standard deviations in case of MENTOR pre-training across all three model architectures. These results indicate that MENTOR provides more efficient use of training data and human annotations, both of which are costly to obtain. Thus, answering RQ2 we conclude that MENTOR increases training efficiency in three domains considered in this work.
6 Conclusion
This paper introduces MENTOR, a novel and intuitively straightforward pre-training method that embeds the “understanding” of domain-specific human salient features into the model, different from previous human perception-guided training approaches, which align the spatial correspondence between human-sourced and model-sourced saliences. To do that, we first trained an autoencoder to learn associations between human-salient features and images of both classes. Once the latent representation between the input image and the human saliency is learned, we used the encoder part to build a domain-specific classifier. We showed that using MENTOR pre-training results in performance gains in classification tasks, even compared to a model whose weights were first transferred from a network trained on a vast number of images (ImageNet). We also demonstrated that MENTOR pre-training achieved better generalization than state-of-the-art approach to incorporate human salience during training via specially-designed loss function. Finally, we show that initialization of classification models with MENTOR weights improves training efficiency by converging slightly faster than models initialized with non-human-perception-driven weights. Future work is dedicated towards exploring the minimum amount and quality (e.g.,experts- vs non-experts-sourced) of human saliency data necessary to benefit from MENTOR pre-training paradigm.